OpenStudy (anonymous):

6the square root of v. − the square root of the quantity 4 times v squared. − the square root of the quantity 36 times v. + the square root of v squared.?

OpenStudy (anonymous):

\[6\sqrt{v}-\sqrt{4v^2}-\sqrt{36v}+\sqrt{v^2}\]

OpenStudy (anonymous):

I need to simplify this

OpenStudy (anonymous):

@phi

OpenStudy (phi):

this is similar to your other problems. have you figured out how to simplify the 2nd term? \[ \sqrt{4v^2} \]

OpenStudy (anonymous):

Expand it and find pairs?

OpenStudy (phi):

yes, can you do that ?

OpenStudy (anonymous):

\[\sqrt{2\times2 \times v \times v}\]

OpenStudy (anonymous):

Pair of twos and pair of v's, then put one on each side?

OpenStudy (phi):

yes, you "pull out a pair" (leaving a 1 in their place)

OpenStudy (anonymous):

\[2v \sqrt{2v+1}?\]

OpenStudy (phi):

\[ \sqrt{2\cdot 2 \cdot v \cdot v} = 2\cdot v\sqrt{1\cdot 1}= 2v\sqrt{1}= 2v \] normally we don't bother with the 1, but it is nice to know it is there. pay attention... you see 2*2 inside the parens. you take them out, and put one 2 outside 2 sqrt(v*v) you see a pair of v's. you pull them out, and put *one* v outside 2*v*sqrt(1) and of course the square root of 1 is 1, and 2*v*1 is just 2*v or 2v

OpenStudy (phi):

you now have \[ 6\sqrt{v}-\sqrt{4v^2}-\sqrt{36v}+\sqrt{v^2} \\6\sqrt{v}-2v-\sqrt{36v}+\sqrt{v^2} \] can you simplify the 3rd term ?

OpenStudy (anonymous):

Put it to the lowest term?

OpenStudy (phi):

Expand it and find pairs

OpenStudy (anonymous):

It's jst 6 and 6

OpenStudy (phi):

it is \[ \sqrt{36v}= \sqrt{6\cdot6\cdot v} \] now pull out any pairs

OpenStudy (anonymous):

\[6\sqrt{v}\]

OpenStudy (phi):

so we now have \[ 6\sqrt{v}-\sqrt{4v^2}-\sqrt{36v}+\sqrt{v^2} \\6\sqrt{v}-2v-\sqrt{36v}+\sqrt{v^2} \\6\sqrt{v}-2v-6\sqrt{v}+\sqrt{v^2}\] now the last term

OpenStudy (anonymous):

\[v \sqrt{v}\]

OpenStudy (phi):

?

OpenStudy (anonymous):

just \[\sqrt{v} ?\]

OpenStudy (phi):

you pull out a pair... if there is nothing left, put in a 1 in the place of the pair. can you do that ?

OpenStudy (anonymous):

\[v \sqrt{1} ??\]

OpenStudy (phi):

yes or just v you now have \[ 6\sqrt{v}-2v-6\sqrt{v}+v\]

OpenStudy (phi):

now combine *like terms* first, re-arrange to \[ 6\sqrt{v}-6\sqrt{v}+v-2v \] can you simplify this ?

OpenStudy (anonymous):

factor out the v's?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

what do you get ?

OpenStudy (anonymous):

I'm a little confused. Put em all to the side?

OpenStudy (anonymous):

\[-12\sqrt{v^2} - 2v\]

OpenStudy (phi):

The first thing is understand what is a *like term* \[ 6\sqrt{v}-6\sqrt{v}+v-2v \] like terms have the same letter to the same power. that means v and -2v are like terms: both have v to the first power 6 sqrt(v) and -6 sqrt(v) are like terms

OpenStudy (phi):

can you combine the like terms 6 sqrt(v) -6 sqrt(v) remember this says you have 6 sqrt(v) take away 6 sqrt(v)

OpenStudy (anonymous):

6sqrt v

OpenStudy (phi):

6 sqrt(v) take away 6 sqrt(v) is ?

OpenStudy (phi):

6 cows take away 6 cows ? 6 apples take away 6 apples ? 6 strawberries minus 6 strawberries ? \[ 6 \sqrt{v} - 6 \sqrt{v} = ?\]

OpenStudy (anonymous):

Oh, sorry. >.< Not thinking. It's 0

OpenStudy (phi):

what about the last two terms v -2v (notice these are *like terms* because they have the same letter )

OpenStudy (anonymous):

2

OpenStudy (phi):

the v's don't disappear v - 2v = 1 v - 2 v = ?

OpenStudy (phi):

1 v take away 2 v's give how many v's ?

OpenStudy (anonymous):

-v

OpenStudy (phi):

so what is the final answer ?

OpenStudy (anonymous):

-v?

OpenStudy (phi):

yes

OpenStudy (phi):

\[ 6\sqrt{v}-\sqrt{4v^2}-\sqrt{36v}+\sqrt{v^2} = -v\]